OpenStudy (anonymous):

Expand and simplify (3a-2)(4a-5)

OpenStudy (anonymous):

(3a - 2)(4a - 5) = 3a(4a - 5) - 2(4a - 5) Can you solve this?

OpenStudy (anonymous):

Nope :(

OpenStudy (anonymous):

Let's start with this. 3a(4a - 5) Can you solve this?

OpenStudy (anonymous):

No the problem is i just really dont get how to do the whole expand and simplify thing :(

OpenStudy (anonymous):

Ok. So, you don't know how to FOIL?

OpenStudy (anonymous):

Noo

OpenStudy (anonymous):

First Outside Inside Last So first, multiply 3a and 4a together.

OpenStudy (anonymous):

Okay so thats 12a?

OpenStudy (anonymous):

That's 12a^2 + a x a = a^2 remember?

OpenStudy (anonymous):

Yeahh

OpenStudy (anonymous):

First: 12a^2 Outside Inside Last Now, multiply 3a and -5

OpenStudy (anonymous):

-15a?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Alright, that's right. First: 12a^2 Outside: -15a Inside Last Now multiply -2 and 4a

OpenStudy (anonymous):

-8a??

OpenStudy (anonymous):

Ok. That's it. First: 12a^2 Outside: -15a Inside: -8a Last Now finally, -2 times -5

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Ok. First: 12a^2 Outside: -15a Inside: -8a Last: 10 Now combine. 12a^2 - 15a - 8a + 10 = 12a^2 - 23a + 10 12a^2 - 23a + 10 is the answer. Do you see hwat terms you have to multiply now? Look at the order I told you to multiply things and where it is in the multiplication. First: 12a^2 Outside: -15a Inside: -8a Last: 10 (3a - 2)(4a - 5)

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

np :]